OpenStudy (anonymous):

(sin^2(x)-cos^2(x))/(sin(x)+cos(x))=(sin(x)-cos(x)) ; prove

OpenStudy (mertsj):

\[\frac{\sin x-\cos x)(\sin x+\cos x)}{\sin x+\cos x}=\sin x-\cos x\]

OpenStudy (mertsj):

Cancel the common factor.